A feline cat isn't showing signs of shock, the cat weighs 8 lbs, and is 10% dehydrated, and doesn't have any vomiting or diarrhea. use this information to calculate the fluid rate roe each needed phase of fluid therapy for this patient. Be sure to include the formula you're using and show each step in the calculation of fluid rates
Answer #1

Calculating the fluid replacement volume ​​​​​:

Body weight in lbs × percent of dehydration (in decimal) × 500 = fluid replacement in ml

Body weight = 8 lbs

Percent of dehydration = 10%

Fluid replacement in ml = 8 × 10 /100 ×500

= 40000 ÷ 100

= 400ml

If the cat had vomiting or diarrhea that fluid loss should be added with this.

Calculating fluid rate according to AAHA :

Fluid rate for cate = 80 × body weight in kg/24 hour

Body weight = 8 lbs

1 kg = 2.205 lbs

8 lbs = 8 ÷ 2.205

= 3.6 kg

Fluid rate = 80 × 3.6

= 288 ml /24 hours

OR

Fluid rate = 2-3 ml/kg/hr as thumb rule

= 2-3 ml×3.6 per hour

= 7.2 - 10.8 ml /hr